# Clarifying Ambiguities: on the Role of Ambiguity Types in Prompting Methods for Clarification Generation

### v1

@@ Abstract.
In information retrieval (IR), providing appropriate clarifications to better understand users’ information needs is crucial for building a proactive search-oriented dialogue system. Due to the strong in-context learning ability of large language models (LLMs), recent studies investigate prompting methods to generate clarifications using few-shot or Chain of Thought (CoT) prompts. However, vanilla CoT prompting does not distinguish the characteristics of different information needs, making it difficult to understand how LLMs resolve ambiguities in user queries. In this work, we focus on the concept of ambiguity for clarification, seeking to model and integrate ambiguities in the clarification process. To this end, we comprehensively study the impact of prompting schemes based on reasoning and ambiguity for clarification. The idea is to enhance the reasoning abilities of LLMs by limiting CoT to predict first ambiguity types that can be interpreted as instructions to clarify, then correspondingly generate clarifications. We name this new prompting scheme Ambiguity Type-Chain of Thought (AT-CoT). Experiments are conducted on various datasets containing human-annotated clarifying questions to compare AT-CoT with multiple baselines. We also perform user simulations to implicitly measure the quality of generated clarifications under various IR scenarios.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italy.booktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italyisbn: 979-8-4007-1592-1/25/07doi: 10.1145/XXXXXX.XXXXXXccs: Information systems Information retrieval
@@ 1. Introduction
To validate the effectiveness of our method, experiments are carried out on both intrinsic and extrinsic tasks (resp. clarification generation and IR) on numerous datasets including Qulac (Aliannejadi et al., 2019), ClariQ (Aliannejadi et al., 2021), and TREC IR collections (Clarke et al., 2009, 2010, 2011, 2012). For the IR task, we follow previous work (Aliannejadi et al., 2019; Zou et al., 2023; ERBACHER et al., 2024) and perform user simulation to generate multi-turn conversations and then transform the generated conversations to reformulated queries for IR. We compare different clarification interaction scenarios such as proposing query reformulations for users to select (select) and asking a single clarifying question for users to respond (respond). To summarize, our main contribution is as follows:
$\bullet$ We analyze ambiguities from the perspective of ambiguity types, comprehensively investigate the impact of integrating
@@ 2.1. Ambiguity in User Queries
Figure 2. Comparison between existing ambiguity type taxonomies for analytical purposes and our proposed action-based one.
@@ 2.2. Clarification in Information Retrieval
Clarification serves the purpose of eliciting the user’s information need (Zamani et al., 2020) by exchanging with the user and exploiting the user’s feedback. The clarification form could be diverse, either by proposing reformulated queries to diversify the retrieval results or by asking clarifying questions to induce users to clarify. Early attempts of clarification generation focus on automatic query expansion (Carpineto and Romano, 2012), whereby users’ original queries are rewritten or augmented. For example, Chirita et al. (2007) proposed to expand user queries with terms collected from user data to handle ambiguities of short keyword queries in web searches. Other studies (Cucerzan and White, 2007; Cao et al., 2008; Mei et al., 2008) investigate query suggestions by exploring different user-specific data sources such as landing pages, clicks, or hitting time. Both query expansion and query suggestions can be regarded as a form of proposing reformulated queries to users and collecting users’ feedback. Recent studies concentrate more on asking clarifying questions (Rao and Daumé III, 2018; Rahmani et al., 2023; Aliannejadi et al., 2019; Xu et al., 2019; Lee et al., 2023). The common approach consists of the conversation history as input to a generative language model to generate CQs.
Before the era of LLM, clarification generation methods mostly consist of training sequence-to-sequence neural models (e.g. seq2seq (Sutskever et al., 2014)) using labeled data. For example, Guo et al. (2021) fine-tune BART (Lewis et al., 2020) to generate clarifying questions provided ambiguous questions as input; Xu et al. (2019) investigated knowledge-based clarifying question generation and concatenated entity texts and the current question as input to a Seq2seq (Bahdanau et al., 2014) model. Recent studies in the post-LLM era have increasingly focused on LLM prompting methods, such as using few-shot prompting (Lee et al., 2023; Zhang et al., 2024), Chain of Thought (CoT) prompting (Deng et al., 2023; Zhang et al., 2024). Our work extends existing studies on LLM prompting methods for clarification generation by integrating ambiguity types into CoT reasoning.
@@ 2.3. Conversation Simulation in Search
In IR, user simulation consists of creating artificial conversations based on hypotheses about user behaviors, often used to automatically test the performance of dialogue systems without performing real user tests (Eckert et al., 1997; Dupret and Piwowarski, 2008; Chuklin et al., 2013). The common approach is to instantiate a user agent to communicate with the dialogue system according to certain strategies (Erbacher et al., 2022; Câmara et al., 2022; Maxwell et al., 2015). Hypotheses about user behavior are made to control how user agents respond, depending on the purpose of the simulation. For example, to test the quality of reformulation in IR systems, Erbacher et al. (2022) assumed that the user agent is greedy and fully cooperative, thus always selecting the reformulations most similar to the user intent. In another work (ERBACHER et al., 2024), user agents are allowed to only respond ’yes’ or ’no’ to augment IR datasets with multi-turn conversations. Some studies involve simulation of more complex user behaviors, in which user agents are initialized with different variables, each corresponding to a specific type of user. Recent studies have increasingly focused on LLM-based conversation simulation. For example, Owoicho et al. (2023) built a user simulator for mixed-initiative multi-turn conversation systems by few-prompting LLMs. Following previous work on LLM-based conversation simulation in IR, in our work, we instantiate our user agent using LLMs and simulate user responses by few-shot LLM prompting.
@@ 3.1. Ambiguity Type Taxonomy
Table 1 presents detailed explanations of our AT taxonomy. To facilitate understanding of the connection between our taxonomy and previous work, we categorize each AT from existing taxonomies within our framework as shown in Figure 2. Compared to previous taxonomies, the strength of our taxonomy lies in its dual function: each AT in our taxonomy not only helps LLMs understand underlying ambiguities, but can also be easily interpreted as an action for LLMs to take.
@@ 3.2. Prompting Formulation for Clarification Generation
This section aims to respond to RQ2, i.e. how to integrate ambiguities, abstracted by the ambiguity type taxonomy in Section 3.1, into reasoning for LLM prompting methods. We aim to achieve this by constraining the reasoning of CoT prompting. Intuitively, we seek to require LLMs to predict ATs in our taxonomy to integrate ambiguities into reasoning, which endows LLMs the capability to reason in a way that we expect and take explainable actions to clarify ambiguous queries. We hypothesize that ambiguity-oriented reasoning is better than freely generated LLM reasoning. Therefore, we propose Ambiguity Type-Chain of Thought (AT-CoT) that extends CoT prompting. To effectively access the impact of integrating ambiguities into LLM reasoning, we use another two prompting schemes as baselines: standard prompting, which simply requires LLMs to generate clarifications without any intermediate steps; AT-standard prompting, for which we add definitions of ATs in our taxonomy into prompt instructions. We use AT-standard to validate the impact of simply informing LLMs of possible ATs without asking LLMs to generate reasoning. Table 2 shows detailed prompt instructions for different prompting schemes. Mathematically, each prompting method can be formulated as follows:
@@ 4.2. Evaluation Protocol
For each query, we generate multiple CQs to fairly evaluate the performance of different prompting methods on the clarification generation (CG) task. Several factors drive this decision: 1) In CG datasets, each query corresponds to numerous human-annotated CQs, covering different clarification possibilities. However, human-annotated CQs cannot contain all possible CQs. Since automatic metrics such as BERTScore (Zhang* et al., 2020) capture the semantic similarity between generated CQs and reference CQs, it is likely that a high-quality generated CQ gets a low BERTScore. To mitigate this issue, we seek to generate multiple diverse CQs, therefore reducing the probability that none of the generated CQs is semantically similar to any of the human-annotated ones. 2) For AT-CoT, since multiple ambiguity types may exist for a query, it is natural to generate multiple CQs to account for different ATs. For other prompting methods that do not predict ATs, generating multiple CQs is also helpful for fair comparison. As evidenced in Wang et al. (2022), the voting strategy can increase the performance of LLM prompting methods. Generating multiple CQs and comparing the best-performing CQ can be regarded as a type of voting, by which we reduce the variance of method performances on CG.
@@
$\bullet$ select: In each turn, 5 RQs are generated. We adopted a moderate temperature of 0.7 to balance the diversity and consistency of the generated RQs, ensuring that they are varied while avoiding excessive creativity. The user agent selects the RQ that best corresponds to their intent, and the conversation continues based on the selected RQ.
$\bullet$ respond: In each turn, a CQ is generated. The user agent responds to it based on the intent of the provided user.
@@
To simulate multi-turn conversations, three prompts are chained: generation, response, reformulation. The generation prompt has four variants, each representing a prompting methods as described in Section 3.2. We use this prompt to generate RQs (scenario select) or CQs (scenario respond). The response prompt generates simulated user responses and has two variants, each corresponding to an interaction scenario. Paragraphs describing user intents and clarifications generated by generation as used as input for the response prompt. Since facet-specific document relevance judgments are provided in TREC Web track 2009-2014, we directly use facets as user intents for response and use gold document relevance labels. For TREC DL Hard, we use each relevant document as user intent for response, then use the same document as the gold label in IR evaluation. For each simulated conversation, the reformulation prompt uses the simulated conversation as input and summarizes the conversation using a reformulated query. The objective of using reformulation is to facilitate the evaluation of IR tasks, since most existing IR models retrieve documents by queries rather than conversations. Examples of the response and reformulation prompts can be found in Table 4. As a result, we have eight types of simulated conversations, each corresponding to a unique (clarification generation prompt, interaction scenario) combination. To simplify the simulation, we assume that users are always cooperative and that no intent shifts occur during the conversation. Each conversation is initialized by a user query and simulated to three turns with no stopping rules. User agents are always provided with complete descriptions of user intents.
Table 4. Prompt examples of respond and reformulation. There is no reformulation prompt for the mode select.
@@
response (scenario select) Imagine that you are a user seeking information with the help of a conversational assistant. At each turn of the conversation, the assistant provides you several reformulated queries to better understand your intention. Given a conversation history and a paragraph describing the user intent, choose the reformulation that most accurately reflects the provided user intent.
@@
response (scenario respond) Imagine that you are a user seeking information with the help of a conversational assistant. At each turn of the conversation, the assistant asks a clarification question to better understand your intention. Given a conversation history and a paragraph describing the user intent, respond to the clarification question based on the provided user intent.
@@
reformulation (scenario respond) Given a conversation history, summarize the conversation as a reformulated query. The chat history includes the initial query and several clarification turns between the user and a virtual assistant.
@@ 4.3. Evaluation Metric
For the IR task, we use different standard metrics following previous work (Clarke et al., 2009; Zhou et al., 2024): We use nDCG@10 (Normalized Discounted Cumulative Gain (Wang et al., 2013)) for TREC web track 2009-2014. For TREC DL Hard, since we use each relevant document as user intent for simulation, then verify whether the target document is ranked higher through clarification, we use MRR@10 (Mean Reciprocal Rank (Radev et al., 2002)) as the evaluation metric.
@@ 4.4. Implementation Details
Following (Nogueira and Cho, 2019; Karpukhin et al., 2020; Nogueira et al., 2020), we adopt a two-stage retriever-rerank pipeline for IR tasks. Top-$K$ relevant documents are first retrieved from a large-scale document collection using BM25 and then reranked using MonoT5. For the retriever, we use a no-tuning pyserini Lucene implementation with $k$ fixed to 100. For the reranker, we use a pre-trained MonoT5 (Nogueira et al., 2020).
@@
We use ir_datasets (MacAvaney et al., 2021), a commonly used Python package in IR community to manage IR datasets. The Python implementation of pyserini [^6] https://github.com/castorini/pyserini is used to build BM25 indexes of Clueweb09 and Clueweb12.
@@ 5. Task 1: Clarification Generation (CG)
To provide more insight into AT-CoT, we analyze the distribution of ambiguity types predicted by AT-CoT. We first investigate the frequency of predicted ATs (namely Semantic, Generalize and Specify), and then focus on the impact of predicting different ATs on the performance of the CG task. Predicted ATs are extracted from the reasoning generated by AT-CoT. Table 7 shows statistics about each group on all CG datasets: the frequency of queries identified as a specific AT and the performance difference in terms of BERTScore between AT-CoT and CoT (in parentheses). Our main conclusions are the following: 1) Semantic and Specify are the most frequent types for all CG datasets, with Specify being slightly more common (i.e. at most 45.9% vs. at most 53.7% for Specify). This observation aligns with the fact that most ATs in existing taxonomies can be categorized as Semantic or Specify. However, though less frequent, the importance of Generalize cannot be overlooked. 2) The Generalize type is more marginal but the fact that 12% of queries in RaoCQ are predicted to be generalizable justifies our decision to include Generalize in our taxonomy. 3) The observation that queries in RaoCQ more often require generalization suggests that the AT predictions of AT-CoT effectively capture the clarification needs of queries and are less likely to be random. Since RaoCQ queries are extracted from user posts on StackExchange, they are generally longer compared to queries in Qulac and ClariQ. It is therefore very likely that a query in RaoCQ does not precisely describe user intents and requires generalization. Differently, queries in Qulac are often short, used for web search, making them less possibly to require generalization. This gap between the frequency of Generalization being predicted and improvements caused by predicting Generalization reflects that AT-CoT adapts well to datasets with different characteristics.

### v2

@@ Abstract.
In information retrieval (IR), providing appropriate clarifications to better understand users’ information needs is crucial for building a proactive search-oriented dialogue system. Due to the strong in-context learning ability of large language models (LLMs), recent studies investigate prompting methods to generate clarifications using few-shot or Chain of Thought (CoT) prompts. However, vanilla CoT prompting does not distinguish the characteristics of different information needs, making it difficult to understand how LLMs resolve ambiguities in user queries. In this work, we focus on the concept of ambiguity for clarification, seeking to model and integrate ambiguities in the clarification process. To this end, we comprehensively study the impact of prompting schemes based on reasoning and ambiguity for clarification. The idea is to enhance the reasoning abilities of LLMs by limiting CoT to predict first ambiguity types that can be interpreted as instructions to clarify, then correspondingly generate clarifications. We name this new prompting scheme Ambiguity Type-Chain of Thought (AT-CoT). Experiments are conducted on various datasets containing humanannotated clarifying questions to compare AT-CoT with multiple baselines. We also perform user simulations to implicitly measure the quality of generated clarifications under various IR scenarios.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3729922isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Information retrieval
@@ 1. Introduction
To validate the effectiveness of our method, experiments are carried out on both intrinsic and extrinsic tasks (resp. clarification generation and IR) on numerous datasets including Qulac (Aliannejadi et al., 2019), ClariQ (Aliannejadi et al., 2021), and TREC IR collections (Clarke et al., 2009, 2010, 2011, 2012). For the IR task, following previous work (Aliannejadi et al., 2019; Zou et al., 2023; ERBACHER et al., 2024), we perform user simulation to generate multi-turn conversations and then transform the generated conversations to reformulated queries. We compare different clarification interaction scenarios such as proposing query reformulations for users to select (select) and asking a single clarifying question for users to respond (respond). To summarize, our main contribution is as follows:
$\bullet$ We analyze ambiguities from the perspective of ambiguity types, comprehensively investigate the impact of integrating ambiguities and reasoning in LLM prompting methods for clarification.
@@ 2.1. Ambiguity in User Queries
Figure 2. Ambiguity type taxonomies for analytical purposes in previous work.
@@ 2.2. Clarification in Information Retrieval
Clarification serves the purpose of eliciting the user’s information need (Zamani et al., 2020) by exchanging with the user and exploiting the user’s feedback. The clarification form could be diverse, either by proposing reformulated queries to diversify the retrieval results or by asking clarifying questions to induce users to clarify. Early attempts of clarification generation focus on automatic query expansion (Carpineto and Romano, 2012), whereby users’ original queries are rewritten or augmented. For example, Chirita et al. (2007) proposed to expand user queries with terms collected from user data to handle ambiguities of short keyword queries in web searches. Other studies (Cucerzan and White, 2007; Cao et al., 2008; Mei et al., 2008) investigate query suggestions by exploring different user-specific data sources such as landing pages, clicks, or hitting time. Both query expansion and query suggestions can be regarded as a form of proposing reformulated queries to users and collecting users’ feedback. Recent studies concentrate more on asking clarifying questions (Rao and Daumé III, 2018; Rahmani et al., 2023; Aliannejadi et al., 2019; Xu et al., 2019; Lee et al., 2023). The common approach consists of using the conversation history as input to a generative language model to generate CQs.
Before the era of LLM, clarification generation methods mostly consist of training sequence-to-sequence neural models (e.g. seq2seq (Sutskever et al., 2014)) using labeled data. For example, Guo et al. (2021) fine-tune BART (Lewis et al., 2020) to generate clarifying questions with ambiguous questions provided as input; Xu et al. (2019) investigated knowledge-based clarifying question generation and concatenated entity texts and the current question as input to a Seq2seq (Bahdanau et al., 2014) model. Recent studies in the post-LLM era have increasingly focused on LLM prompting methods, such as using few-shot prompting (Lee et al., 2023; Zhang et al., 2024), Chain of Thought (CoT) prompting (Deng et al., 2023; Zhang et al., 2024). Our work extends existing studies on LLM prompting methods for clarification generation by integrating ambiguity types into CoT reasoning.
@@ 2.3. Conversation Simulation in Search
In IR, user simulation consists of creating artificial conversations based on hypotheses about user behaviors, often used to automatically test the performance of dialogue systems without performing real user tests (Eckert et al., 1997; Dupret and Piwowarski, 2008; Chuklin et al., 2013). The common approach is to instantiate a user agent to communicate with the dialogue system according to certain strategies (Erbacher et al., 2022; Câmara et al., 2022; Maxwell et al., 2015). Hypotheses about user behavior are made to control how user agents respond, depending on the purpose of the simulation. For example, to test the quality of reformulation in IR systems, Erbacher et al. (2022) assumed that the user agent is greedy and fully cooperative, thus always selecting the reformulations most similar to the user intent. In another work (ERBACHER et al., 2024), user agents are allowed to only respond ’yes’ or ’no’ to augment IR datasets with multi-turn conversations. Some studies involve simulation of more complex user behaviors, in which user agents are initialized with different variables, each corresponding to a specific type of user. Recent studies have increasingly focused on LLM-based conversation simulation. For example, Owoicho et al. (2023) built a user simulator for mixed-initiative multi-turn conversation systems by prompting LLMs. Following previous work on LLM-based conversation simulation in IR, in our work, we instantiate our user agent using LLMs and simulate user responses by few-shot LLM prompting.
@@ 3.1. Ambiguity Type Taxonomy
Table 1 presents detailed explanations of our AT taxonomy. Compared to previous taxonomies, the strength of our taxonomy lies in its dual function: each AT in our taxonomy not only helps LLMs understand underlying ambiguities, but can also be easily interpreted as an action for LLMs to take.
@@ 3.2. Prompting Formulation for Clarification Generation
This section aims to respond to RQ2, i.e. how to integrate ambiguities, abstracted by the ambiguity type taxonomy in Section 3.1, into reasoning for LLM prompting methods. We aim to achieve this by constraining the reasoning of CoT prompting. Intuitively, we seek to require LLMs to predict ATs in our taxonomy to integrate ambiguities into reasoning, which endows LLMs the capability to reason in a way that we expect and take explainable actions to clarify ambiguous queries. We hypothesize that ambiguity-oriented reasoning is better than freely generated LLM reasoning. Therefore, we propose Ambiguity Type-Chain of Thought (AT-CoT) that extends CoT prompting. To effectively access the impact of integrating ambiguities into LLM reasoning, we use another two prompting schemes as baselines: standard prompting, which simply requires LLMs to generate clarifications without any intermediate steps; AT-standard prompting, for which we add definitions of ATs in our taxonomy into prompt instructions. We use AT-standard to validate the impact of simply informing LLMs of possible ATs without asking LLMs to generate reasoning. Table 2 shows detailed system instructions for different prompting schemes. Mathematically, each prompting method can be formulated as follows:
@@ 4.2. Evaluation Protocol
For each query, we generate multiple CQs to fairly evaluate the performance of different prompting methods on the clarification generation (CG) task. Several factors drive this decision: 1) In CG datasets, each query corresponds to numerous human-annotated CQs, covering different clarification possibilities. However, human-annotated CQs cannot contain all possible CQs. Since automatic metrics such as BERTScore (Zhang* et al., 2020) capture the semantic similarity between generated CQs and reference CQs, it is likely that a high-quality generated CQ gets a low BERTScore. To mitigate this issue, we seek to generate multiple diverse CQs, therefore reducing the probability that none of the generated CQs is semantically similar to any of the human-annotated ones. 2) For AT-CoT, since multiple ambiguity types may exist for a query, it is natural to generate multiple CQs to account for different ATs. For other prompting methods that do not predict ATs, generating multiple CQs is also helpful for fair comparison. As evidenced in Wang et al. (2022), the voting strategy can increase the performance of LLM prompting methods. Generating multiple CQs and comparing the best-performing CQ can be regarded as a type of voting, by which we reduce the variance of prompting performances on CG.
@@
$\bullet$ select: In each turn, 5 RQs are generated. We adopted a moderate temperature of 0.6 to balance the diversity and consistency of the generated RQs, ensuring that they are varied while avoiding excessive creativity. The user agent selects the RQ that best corresponds to their intent, and the conversation continues based on the selected RQ.
$\bullet$ respond: In each turn, a CQ is generated. The user agent responds to it based on the provided user intent.
@@
To simulate multi-turn conversations, three prompts are chained: generation, response, reformulation. The generation prompt has four variants, each representing a prompting methods as described in Section 3.2. We use this prompt to generate RQs (scenario select) or CQs (scenario respond). The response prompt generates simulated user responses and has two variants, each corresponding to an interaction scenario. The clarifications generated by generation and paragraphs describing user intents are used as input for the response prompt. Since facet-specific document relevance judgments are provided in TREC Web track 2009-2014, we directly use facets as user intents for response and use gold document relevance labels. For TREC DL Hard, we use each relevant document as user intent for response, then use the same document as the gold label in IR evaluation. For each simulated conversation, the reformulation prompt uses the simulated conversation as input and summarizes the conversation into a reformulated query. The objective of using reformulation is to facilitate the evaluation of IR tasks, since most existing IR models retrieve documents by queries rather than conversations. Examples of the response and reformulation prompts can be found in Table 4. As a result, we have eight types of simulated conversations, each corresponding to a unique (clarification generation prompt, interaction scenario) combination. To simplify the simulation, we assume that users are always cooperative and that no intent shifts occur during the conversation. Each conversation is initialized by a user query and simulated to three turns with no stopping rules. User agents are always provided with complete descriptions of user intents.
Table 4. Prompts of respond and reformulation. There is no reformulation prompt for the mode select.
@@
response (scenario select) Imagine that you are a user seeking information with the help of a conversational assistant. At each turn of the conversation, the assistant provides you several reformulated queries to better understand your intent. Given a conversation history and a paragraph describing the user intent, choose the reformulated query that most accurately reflects the provided user intent.
@@
response (scenario respond) Imagine that you are a user seeking information with the help of a conversational assistant. At each turn of the conversation, the assistant asks a clarification question to better understand your intent. Given a conversation history and a paragraph describing the user intent, respond to the clarification question based on the provided user intent.
@@
reformulation (scenario respond) Given a conversation history, summarize the conversation as a reformulated query. The conversation history includes the initial query and several clarification turns between the user and a virtual assistant.
@@ 4.3. Evaluation Metric
For the IR task, we use different standard metrics following previous work (Clarke et al., 2009; Zhou et al., 2024): We use nDCG@10 (Normalized Discounted Cumulative Gain (Wang et al., 2013)) for TREC web track 2009-2014. For TREC DL Hard, since we use each relevant document as user intent for simulation then verify whether the target document is ranked higher through clarification, we use MRR@10 (Mean Reciprocal Rank (Radev et al., 2002)) as the evaluation metric.
@@ 4.4. Implementation Details
Following (Nogueira and Cho, 2019; Karpukhin et al., 2020; Nogueira et al., 2020), we adopt a two-stage retriever-rerank pipeline for IR tasks. Top-$k$ relevant documents are first retrieved from a large-scale document collection using BM25 and then reranked using MonoT5. For the retriever, we use a no-tuning pyserini [^6] https://github.com/castorini/pyserini Lucene implementation with $k$ fixed to 100. For the reranker, we use a pre-trained MonoT5 (Nogueira et al., 2020).
@@
We use ir_datasets (MacAvaney et al., 2021), a commonly used Python package in IR community to manage IR datasets. The Python implementation of pyserini is used to build BM25 indexes of Clueweb09 and Clueweb12.
@@ 5. Task 1: Clarification Generation (CG)
To provide more insight into AT-CoT, we analyze the distribution of ambiguity types predicted by AT-CoT. We first investigate the frequency of predicted ATs (namely Semantic, Generalize and Specify), and then focus on the impact of predicting different ATs on the performance of the CG task. Predicted ATs are extracted from the reasoning generated by AT-CoT. Table 7 shows statistics about each group on all CG datasets: the frequency of queries identified as a specific AT and the performance difference in terms of BERTScore between AT-CoT and CoT (in parentheses). Our main conclusions are the following: 1) Semantic and Specify are the most frequent types for all CG datasets, with Specify being slightly more common (i.e. at most 45.9% vs. at most 53.7% for Specify). This observation aligns with the fact that most ATs in existing taxonomies can be categorized as Semantic or Specify. However, though less frequent, the importance of Generalize cannot be overlooked. 2) The Generalize type is more marginal but the fact that 12% of queries in RaoCQ are predicted to be generalizable justifies our decision to include Generalize in our taxonomy. 3) The observation that queries in RaoCQ more often require generalization suggests that the AT predictions of AT-CoT effectively capture the clarification needs of queries and are less likely to be random. Since RaoCQ queries are extracted from user posts on StackExchange, they are generally longer compared to queries in Qulac and ClariQ. It is therefore very likely that a query in RaoCQ does not precisely describe user intents and requires generalization. Differently, queries in Qulac are often short, used for web search, making them less possibly to require generalization. This gap between the frequency of Generalize being predicted and improvements caused by predicting Generalize reflects that AT-CoT adapts well to datasets with different characteristics.